Please Answer all parts

Financial data reported by Golf Inc. for the 4th quarter of 2017 is shown below. Note a quarter is 13 weeks.

Net Revenue

Cost of Sales

Inventory (Avg)

Total Assets

Total Loans

870,000

665,000

230,000

1,150,000

290,000

  1. What is Golf’s inventory turnover?   
  2. What is Golf’s percent assets committed to inventory expressed in decimal fraction format?   
  3. How many weeks of inventory are held by Golf?   
Answer #1

(a)

Inventory turnover= Cost of Goods Sold/ Average Inventory

                                = 665000/ 230000= 2.89

So 2.89 times the average inventory is sold during an year for Golf Inc. This shows, the sales is good year around. Ratio can vary, depending on inventory item and size of the company.

(b) Percent assets committed to inventory= Value of inventory/Value of total assets

                                                                = 230000/ 1150000 = 0.2

It shows how good is Golf Inc.’s inventory efficiency. The lower it is, better the profitability can be. It is very critical ratio, especially in case of retail products or during any seasonal demands to maintain the bottom line of a company.

(c)

Weeks of inventory are held:

This shows number of weeks a company will take to sell its average inventory. It shows how quickly company can rotate or needs to rotate the inventory

= Number of weeks per year/ Inventory turnover ratio

= 52/ 2.89= 17.99 weeks